OpenStudy (anonymous):

Solve by finding the value of y in the 1st equation. 3x + y = 10 x - y = 2.

OpenStudy (turingtest):

Not sure about the phrasing here, there are many ways to find y in the first eqn. 3x + y = 10 (1) x - y = 2 (2) solving eqn (1) for y gives y=10-3x sub into (2): x-(10-3x)=2 4x-10=2 4x=12 x=3 back to our expression for y: y=10-3x=10-3(3)=10-9=1 x=3 y=1

OpenStudy (anonymous):

3x + y = 10 ----- (i) y = 10 - 3x Substituting this value of y in (ii) x - y = 2. ----- (ii) x - (10 - 3x) = 2 x - 10 + 3x = 2 4x = 2 + 10 x = 12/4 x = 3 now substitute this value of x in (i) 3(3) + y = 10 y = 10 - 9 y = 1 so solution set (x,y) = (3,1)